OpenStudy (anonymous):

another representation in polar coordinates, of the point ( 1, 5pi/6) is (____, 11pi/6)

ganeshie8 (ganeshie8):

hint : \(x = r \cos(\theta) = -r \cos(\theta + \pi)\)

OpenStudy (anonymous):

lol it would be if i knew what to do. i don't know anything.

ganeshie8 (ganeshie8):

hahah ikr... that was a very cryptic hint lol

OpenStudy (anonymous):

lol... sorry i suck at this

ganeshie8 (ganeshie8):

okay here a more useful hint : \((r, \theta ) = (-r, ~\theta + \pi)\)

OpenStudy (anonymous):

lol.. i just said idk ANYTHING. seriously

OpenStudy (anonymous):

i'm failing pre-cal with a 31

ganeshie8 (ganeshie8):

u wont fail... start somewhere and study module by module lol

ganeshie8 (ganeshie8):

if u want, the answer is `-1`

ganeshie8 (ganeshie8):

( 1, 5pi/6) is same as (-1, 11pi/6)

OpenStudy (anonymous):

i have two pre-cal classes.. for i failed the one in first semester.. its frustrating.

ganeshie8 (ganeshie8):

precalc is very imp if u wanto pursue math career... doing it two times makes u strong only :)

ganeshie8 (ganeshie8):

wish u good luck !